Im not sure what consecutive integers are.
Question- if three times the smaller of two consecutive integers is added to four times the larger, the result is 137. Find the integers?

Answer by sudhanshu_kmr(1152)


two consecutive integers are x and x+1.
3x + 4(x+1) = 137
=> 7x = 133
=> x= 133/7 = 19
so, integers are 19 and 20.